LAMpoly status Sept 10,2002
booth at snf.stanford.edu
Tue Sep 10 14:39:17 PDT 2002
A new hard disk & disk controller have been installed
& checked to verify that the machine will indeed boot
up. I'm scheduled to pick up a new floppy drive
tomorrow morning & install that. If everything is
working, I should be able to install the old recipes
onto the new hard drive, & test to make sure the
recipes run as before. The etch chamber will be wet
cleaned, since it has been turned off for a while.
So we should be running normally by Thursday.
More information about the lampoly